# Advancing Natural Language Formalization to First Order Logic with Fine-tuned LLMs

### v1

@@ 2 Related Work
Translating NL to FOL can be used to enable logical reasoning with natural language. Early approaches to this problem relied mainly on symbolic methods. They used rule-based systems with hand-crafted linguistic rules and syntactic patterns, maping natural language structures directly to logical expressions. Pease and Murray developed an early translator from controlled English to logic for ontology-based knowledge representation [25]. Bos and Markert explored logical inference techniques for recognizing textual entailment [2]. They used model building from automated reasoning to approximate entailment relationships. Other systems often utilize frameworks like Combinatory Categorial Grammar (CCG) to build compositional semantic representations which can then be transformed to lambda expressions, FOL, or other formal languages. For example, ccg2lambda is a compositional semantics system generating logical semantic representations by combining semantic formulas bottom-up, guided by CCG parse trees [19].
Barker-Plummer et al. systematically analysed the difficulties of students during a translation task. They identified, among other, the ambiguity of natural language, complex sentence structures, and uninformative automated feedback as the students key challenges in mapping natural language to formal logic [1]. Cai and Yates addressed the identified scalability by developing techniques for large-scale semantic parsing [4]. They used schema matching and lexicon extension to improve performance on database-grounded tasks. Martínez-Gómez et al. further refined compositional semantics with the ccg2lambda system [19]. This system produces logical semantic representations by composing formulas bottom-up from CCG parse trees.
More recently, researchers have explored methods beyond purely symbolic systems. Lu et al. demonstrated the use of Dual Reinforcement Learning (DRL) to translate NL to both Propositional and First-Order Logic [18].
@@
LogicLLaMA, a LLaMA-7B model fine-tuned specifically for NL-FOL translation, achieved performance comparable to GPT-4, but at a much lower computational cost [30]. During their work they created the dataset MALLS, a dataset of 34K high-quality NL-FOL pairs generated using GPT-4. Other frameworks combine LLMs with symbolic reasoners. For example, LogicLM first translates NL to FOL using LLMs and then applies symbolic reasoning [24]. Experiments showed this combination performed better than using the LLM directly for reasoning. Building on top of that, Lalwani et al. presented NL2FOL [12], a framework to translate NL to FOL in multiple steps with few-shot prompting of an LLM. It is designed to detect logical fallacies in natural language statements.
Thatikonda et al. focused on fine-tuning LLMs for NL to FOL translation [29]. They explored various fine-tuning strategies and data augmentation techniques. Their results show that fine-tuning significantly improves performance compared to zero-shot settings. They also found that incrementally translating premises and conclusions of (natural language) logical arguments can improve logical inference checking.
@@
For the evaluation of NL to FOL translation tasks, various metrics are used in the literature. A common method is an exact, character-by-character match to evaluate whether a prediction of a model is correct [14, 18]. However, this metric is very strict as it does not account for syntactic variations in FOL expressions which do not change the underlying semantics.
Other approaches use logical equivalence as a more robust metric. Using automated theorem provers this metric checks whether the predicted FOL expression is logically equivalent to the ground truth expression, even if they are syntactically different [3, 10]. However, this approach either requires static signatures ensuring that the same predicates are used in both expressions or a normalization step to align predicate names. Experiments which include NL to FOL translation only as a subtask of logical entailment use the accuracy of the entailment task as a metric [29, 12].
Yang et al. defined a new metric called “Logical Equivalence” which is also used in other experiments [30, 16]. In contrast to the previous methods, which rate a prediction either “correct” or “incorrect”, this metric assigns a score based on the number of matching propositional valuations. However, it is questionable whether this metric is suitable for FOL, as it does not account for the complexity of quantifiers and a prediction which differs in only one valuation can still be worse than a prediction which differs in many valuations. For example, $A\wedge B\wedge C\wedge D$ is closer to $A\wedge B$ than to $\mathit{false}$, but differs from the former for $\frac{3}{16}$ of the valuations, but from latter only for $\frac{1}{16}$.
@@ 3.1 Model Selection
For our formalization task, we employ both encoder-decoder and decoder-only large language models (LLMs) with different sizes to enable a systematic comparison across architectures and model sizes. Encoder-decoder models, such as T5-base [26], T5-3B [26], and Flan-T5-XXL [5], are chosen based on their successful application to structured prediction tasks like text-to-text translation and semantic parsing [15], which parallel the formality and compositionality of natural language to logic translation. For the decoder-only setting, we fine-tune LLaMA3.1-8B [9], which is widely used in recent research on fine tuning LLMs for formal language and reasoning tasks [30], as well as Mistral-24B, recognized for its strong performance and efficiency [20]. Additionally, we include Olmo-32B [22] as a representative of modern open-source approaches in large-scale language modeling.
@@
T5-base Encoder-Decoder 220M [26]
T5-3B Encoder-Decoder 3B [26]
Flan-T5-XXL Encoder-Decoder 11B [5]
Olmo-32B Decoder-Only 32B [22]
Mistral-24B Decoder-Only 24B [20]
LLaMA3.1-8B Decoder-Only 8B [9]
@@ 3.2 Datasets
For our experiments, we leverage two of the largest publicly available Natural Language to First-Order Logic (NL-FOL) datasets: MALLS [30] and Willow [8]. These datasets were selected because their considerable size makes them more suitable for training large models than, for instance, FOLIO [10], and they contain significantly more complex natural language utterances compared to simpler resources such as Text2Log [14].
@@ 3.3 Fine-tuning Procedure
To adapt the selected pre-trained language models to the NL-to-FOL formalization task, we fine-tune each model on the combined training split described above. For the decoder-only models (LLaMA3.1-8B, Mistral-24B, and Olmo-32B), we apply Low-Rank Adaptation (LoRA) [11] with rank $r=16$, scaling factor $\alpha=32$, a dropout rate of 0.05, and targeting all attention and feed-forward projection layers, following recommended settings for large-scale language modeling [11]. For Flan-T5-XXL, LoRA was analogously applied to attention and feed-forward layers. T5-base and T5-3B were trained with full fine-tuning. In all LoRa settings mixed precision training was enabled.
All models were trained with the AdamW optimizer [17], a maximum of 12 epochs, early stopping (patience 4) based on validation loss, and cosine (Decoder-Only models)/linear (T5 models) learning rate scheduling with a warmup ratio of 0.05 and weight decay of 0.01. Batch sizes and learning rates are detailed in Table 3. Multi-GPU training was used for larger models as required.
@@ 8. Comparison to Other Approaches
While previous experiments evaluated the performance of our models either internally or in comparison to other neural methods following the same approach, this section extends the analysis by comparing LLaMA3.1-8B to LLaMA-2 13B [29] and the symbolic system ccg2lambda. For the comparison with LLaMA-2 13B, we use the eval subset of the FOLIO dataset, noticing that our model has not been trained on FOLIO. For the evaluation against ccg2lambda, we preprocessed the natural language inputs by removing all commas and appending a period to each sentence, in accordance with the input format expected by the tool.
@@ 3.5 Evaluation
To account for syntactically different but semantically equivalent expressions, we determine logical equivalence using an automated theorem proving (ATP) approach. Both predicted and ground truth FOL expressions are parsed using a context-free grammar that we specifically designed and implemented with the Lark parsing library [27]. The resulting syntax trees are then transformed into Z3 format. To check for logical equivalence, we construct a Z3 solver [21] to test the unsatisfiability of the negated bi-implication $\lnot(\varphi\leftrightarrow\psi)$, where $\varphi$ denotes the model output and $\psi$ the ground truth formula. We use the Z3 solver with a timeout of 10 seconds for each check. If the solver establishes unsatisfiability within this timeframe, the formulas are considered logically equivalent. Our approach follows the general structure proposed by [3].
@@
To further disentangle errors caused by incorrect predicate naming from structural or logical errors, we introduce a predicate matching step. For each predicted FOL expression, predicates are mapped to ground truth predicates using the normalized Levenshtein distance (threshold 0.6) [13]. Once predicates are aligned, both exact match and logical equivalence are recomputed on the normalized formulas. This procedure provides additional insight into whether errors are due to predicate identification or deeper logical misunderstandings.
@@
For Experiment 1, we first establish a baseline by evaluating each model on the dataset without any fine-tuning, in order to verify that fine-tuning indeed improves performance. We further compare the fine-tuned models to strong proprietary, open-source and reasoning models—namely, GPT-4o [23], DeepSeek-R1-0528 [7] with a hybrid 2500 tokens CoT setting, DeepSeek-V3-0324 [6] and LogicLLaMA [30]—for an additional performance reference. Notably, LogicLLaMA was trained on the MALLS dataset but not on Willow, providing an informative benchmark for both in-domain and out-of-domain generalization. In addition, we evaluate the symbolic system ccg2lambda on the dataset to obtain a symbolic reference point for the neural baselines established in this experiment.
@@ 4.3 Results for logical arguments
[29] argue that with traditional datasets, it is difficult to assess the quality of NL to FOL translations automatically. As an automated santiy check, they suggest to use hand-crafted sets of natural language logical arguments, whose (in)validity is known. Then, after translation to FOL, (in)validity of the argument can be checked with an automated theorem prover, and if the validity disagrees with the manually annoteted one, an error in the translation has been spotted. FOLIO [10] is such a dataset of NL arguments.
Our model achieves competitive results (39.70% valid conclusions) compared to the best version of LLaMA-2 13B (37.44% valid conlsusions with inkremental fine-tuning and predicate verifier), even though it was not trained on the FOLIO dataset. This suggests that the iterative fine-tuning procedure on FOLIO proposed by previous work [29] is not strictly required to attain strong performance. Additionally, our model exhibits robust behavior in multi-premise inference settings, where earlier premises are provided as contextual input—despite such scenarios not being part of the training data. This points to the model’s capacity to generalize to more complex reasoning setups without explicit supervision.
@@ 5 Conclusion
For the task of translating natural language into first-order logic, we have fined-tuned several large language models using the MALLS and Willow datasets, and have conducted experiments with several settings. An important aspect was the provision or learning of a list of predicate names that shall occur in the translation. While the provision of ground-truth predicate names greatly helps, learning them in a two-step process did not lead to good results. Our fine-tuned Flan-T5-XXL model with 11 billion parameters, can outperform baselines like GPT-4o, LogicLLaMA and even DeepSeek-R1-0528, as well as a symbolic baseline. The accuracy of translations of natural language arguments from the FOLIO dataset is comparable to the baseline (note that here only (in)validity of the translated argument is considered). The source code for our fine-tuned models is available on GitHub (link), and the models themselves can be accessed on Hugging Face (link).
A common mistake in NL to FOL translation that is also present in the MALLS dataset is the confusion of hard-to-recognise Aristotelian forms: “A pineapple is a fruit with spiky skin” is translated to
@@
We observe that reasoning strategies such as CoT can sometimes degrade performance in NL-to-FOL translation, as initially correct outputs are overwritten during the reasoning process. This aligns with the illusion of thinking phenomenon described by Apple [28]. Future work may benefit from a two-stage approach, where base model outputs are first generated and then verified or refined using a reasoning model, as suggested by Thatikonda et al [29].

### v2

@@ 2 Related Work
Translating NL to FOL can be used to enable logical reasoning with natural language. Early approaches to this problem relied mainly on symbolic methods. They used rule-based systems with hand-crafted linguistic rules and syntactic patterns, maping natural language structures directly to logical expressions. Pease and Murray developed an early translator from controlled English to logic for ontology-based knowledge representation [pease2003english]. Bos and Markert explored logical inference techniques for recognizing textual entailment [bos2005recognising]. They used model building from automated reasoning to approximate entailment relationships. Other systems often utilize frameworks like Combinatory Categorial Grammar (CCG) to build compositional semantic representations which can then be transformed to lambda expressions, FOL, or other formal languages. For example, ccg2lambda is a compositional semantics system generating logical semantic representations by combining semantic formulas bottom-up, guided by CCG parse trees [martinez2016ccg2lambda].
Barker-Plummer et al. systematically analysed the difficulties of students during a translation task. They identified, among other, the ambiguity of natural language, complex sentence structures, and uninformative automated feedback as the students key challenges in mapping natural language to formal logic [barker2009dimensions]. Cai and Yates addressed the identified scalability by developing techniques for large-scale semantic parsing [cai2013LargescaleSemanticParsing]. They used schema matching and lexicon extension to improve performance on database-grounded tasks. Martínez-Gómez et al. further refined compositional semantics with the ccg2lambda system [martinez2016ccg2lambda]. This system produces logical semantic representations by composing formulas bottom-up from CCG parse trees.
More recently, researchers have explored methods beyond purely symbolic systems. Lu et al. demonstrated the use of Dual Reinforcement Learning (DRL) to translate NL to both Propositional and First-Order Logic [lu2022ParsingNaturalLanguage].
@@
LogicLLaMA, a LLaMA-7B model fine-tuned specifically for NL-FOL translation, achieved performance comparable to GPT-4, but at a much lower computational cost [yang2024HarnessingPowerLarge]. During their work they created the dataset MALLS, a dataset of 34K high-quality NL-FOL pairs generated using GPT-4. Other frameworks combine LLMs with symbolic reasoners. For example, LogicLM first translates NL to FOL using LLMs and then applies symbolic reasoning [pan2023logiclmempoweringlargelanguage]. Experiments showed this combination performed better than using the LLM directly for reasoning. Building on top of that, Lalwani et al. presented NL2FOL [lalwani2024nl2fol], a framework to translate NL to FOL in multiple steps with few-shot prompting of an LLM. It is designed to detect logical fallacies in natural language statements.
Thatikonda et al. focused on fine-tuning LLMs for NL to FOL translation [thatikonda2024strategies]. They explored various fine-tuning strategies and data augmentation techniques. Their results show that fine-tuning significantly improves performance compared to zero-shot settings. They also found that incrementally translating premises and conclusions of (natural language) logical arguments can improve logical inference checking.
@@
For the evaluation of NL to FOL translation tasks, various metrics are used in the literature. A common method is an exact, character-by-character match to evaluate whether a prediction of a model is correct [levkovskyi2021GeneratingPredicateLogic, lu2022ParsingNaturalLanguage]. However, this metric is very strict as it does not account for syntactic variations in FOL expressions which do not change the underlying semantics.
Other approaches use logical equivalence as a more robust metric. Using automated theorem provers this metric checks whether the predicted FOL expression is logically equivalent to the ground truth expression, even if they are syntactically different [brunello2025EvaluatingLLMsCapabilities, han2022folio]. However, this approach either requires static signatures ensuring that the same predicates are used in both expressions or a normalization step to align predicate names. Experiments which include NL to FOL translation only as a subtask of logical entailment use the accuracy of the entailment task as a metric [thatikonda2024strategies, lalwani2024nl2fol].
Yang et al. defined a new metric called “Logical Equivalence” which is also used in other experiments [yang2024HarnessingPowerLarge, liu2025few]. In contrast to the previous methods, which rate a prediction either “correct” or “incorrect”, this metric assigns a score based on the number of matching propositional valuations. However, it is questionable whether this metric is suitable for FOL, as it does not account for the complexity of quantifiers and a prediction which differs in only one valuation can still be worse than a prediction which differs in many valuations. For example, $A\wedge B\wedge C\wedge D$ is closer to $A\wedge B$ than to $\mathit{false}$, but differs from the former for $\frac{3}{16}$ of the valuations, but from latter only for $\frac{1}{16}$.
@@ 3.1 Model Selection
For our formalization task, we employ both encoder-decoder and decoder-only large language models (LLMs) with different sizes to enable a systematic comparison across architectures and model sizes. Encoder-decoder models, such as T5-base [raffel2020exploring], T5-3B [raffel2020exploring], and Flan-T5-XXL [chung2024scaling], are chosen based on their successful application to structured prediction tasks like text-to-text translation and semantic parsing [li2023t5], which parallel the formality and compositionality of natural language to logic translation. For the decoder-only setting, we fine-tune LLaMA3.1-8B [grattafiori2024llama], which is widely used in recent research on fine tuning LLMs for formal language and reasoning tasks [yang2024HarnessingPowerLarge], as well as Mistral-24B, recognized for its strong performance and efficiency [mistral2025small3]. Additionally, we include Olmo-32B [olmo20242] as a representative of modern open-source approaches in large-scale language modeling.
@@
T5-base Encoder-Decoder 220M [raffel2020exploring]
T5-3B Encoder-Decoder 3B [raffel2020exploring]
Flan-T5-XXL Encoder-Decoder 11B [chung2024scaling]
Olmo-32B Decoder-Only 32B [olmo20242]
Mistral-24B Decoder-Only 24B [mistral2025small3]
LLaMA3.1-8B Decoder-Only 8B [grattafiori2024llama]
@@ 3.2 Datasets
For our experiments, we leverage two of the largest publicly available Natural Language to First-Order Logic (NL-FOL) datasets: MALLS [yang2024HarnessingPowerLarge] and Willow [deveci2024transformer]. These datasets were selected because their considerable size makes them more suitable for training large models than, for instance, FOLIO [han2022folio], and they contain significantly more complex natural language utterances compared to simpler resources such as Text2Log [levkovskyi2021GeneratingPredicateLogic].
@@ 3.3 Fine-tuning Procedure
To adapt the selected pre-trained language models to the NL-to-FOL formalization task, we fine-tune each model on the combined training split described above. For the decoder-only models (LLaMA3.1-8B, Mistral-24B, and Olmo-32B), we apply Low-Rank Adaptation (LoRA) [hu2022lora] with rank $r=16$, scaling factor $\alpha=32$, a dropout rate of 0.05, and targeting all attention and feed-forward projection layers, following recommended settings for large-scale language modeling [hu2022lora]. For Flan-T5-XXL, LoRA was analogously applied to attention and feed-forward layers. T5-base and T5-3B were trained with full fine-tuning. In all LoRa settings mixed precision training was enabled.
All models were trained with the AdamW optimizer [loshchilov2017decoupled], a maximum of 12 epochs, early stopping (patience 4) based on validation loss, and cosine (Decoder-Only models)/linear (T5 models) learning rate scheduling with a warmup ratio of 0.05 and weight decay of 0.01. Batch sizes and learning rates are detailed in Table 3. Multi-GPU training was used for larger models as required.
@@ 8. Comparison to Other Approaches
While previous experiments evaluated the performance of our models either internally or in comparison to other neural methods following the same approach, this section extends the analysis by comparing LLaMA3.1-8B to LLaMA-2 13B [thatikonda2024strategies] and the symbolic system ccg2lambda. For the comparison with LLaMA-2 13B, we use the eval subset of the FOLIO dataset, noticing that our model has not been trained on FOLIO. For the evaluation against ccg2lambda, we preprocessed the natural language inputs by removing all commas and appending a period to each sentence, in accordance with the input format expected by the tool.
@@ 3.5 Evaluation
To account for syntactically different but semantically equivalent expressions, we determine logical equivalence using an automated theorem proving (ATP) approach. Both predicted and ground truth FOL expressions are parsed using a context-free grammar that we specifically designed and implemented with the Lark parsing library [larkparser2024]. The resulting syntax trees are then transformed into Z3 format. To check for logical equivalence, we construct a Z3 solver [z3prover2024] to test the unsatisfiability of the negated bi-implication $\lnot(\varphi\leftrightarrow\psi)$, where $\varphi$ denotes the model output and $\psi$ the ground truth formula. We use the Z3 solver with a timeout of 10 seconds for each check. If the solver establishes unsatisfiability within this timeframe, the formulas are considered logically equivalent. Our approach follows the general structure proposed by [brunello2025EvaluatingLLMsCapabilities].
@@
To further disentangle errors caused by incorrect predicate naming from structural or logical errors, we introduce a predicate matching step. For each predicted FOL expression, predicates are mapped to ground truth predicates using the normalized Levenshtein distance (threshold 0.6) [levenshtein1966binary]. Once predicates are aligned, both exact match and logical equivalence are recomputed on the normalized formulas. This procedure provides additional insight into whether errors are due to predicate identification or deeper logical misunderstandings.
@@
For Experiment 1, we first establish a baseline by evaluating each model on the dataset without any fine-tuning, in order to verify that fine-tuning indeed improves performance. We further compare the fine-tuned models to strong proprietary, open-source and reasoning models—namely, GPT-4o [openai2024gpt4o], DeepSeek-R1-0528 [deepseekai2025deepseekr1incentivizingreasoningcapability] with a hybrid 2500 tokens CoT setting, DeepSeek-V3-0324 [deepseekai2024deepseekv3technicalreport] and LogicLLaMA [yang2024HarnessingPowerLarge]—for an additional performance reference. Notably, LogicLLaMA was trained on the MALLS dataset but not on Willow, providing an informative benchmark for both in-domain and out-of-domain generalization. In addition, we evaluate the symbolic system ccg2lambda on the dataset to obtain a symbolic reference point for the neural baselines established in this experiment.
@@ 4.3 Results for logical arguments
[thatikonda2024strategies] argue that with traditional datasets, it is difficult to assess the quality of NL to FOL translations automatically. As an automated santiy check, they suggest to use hand-crafted sets of natural language logical arguments, whose (in)validity is known. Then, after translation to FOL, (in)validity of the argument can be checked with an automated theorem prover, and if the validity disagrees with the manually annoteted one, an error in the translation has been spotted. FOLIO [han2022folio] is such a dataset of NL arguments.
Our model achieves competitive results (39.70% valid conclusions) compared to the best version of LLaMA-2 13B (37.44% valid conlsusions with inkremental fine-tuning and predicate verifier), even though it was not trained on the FOLIO dataset. This suggests that the iterative fine-tuning procedure on FOLIO proposed by previous work [thatikonda2024strategies] is not strictly required to attain strong performance. Additionally, our model exhibits robust behavior in multi-premise inference settings, where earlier premises are provided as contextual input—despite such scenarios not being part of the training data. This points to the model’s capacity to generalize to more complex reasoning setups without explicit supervision.
@@ 5 Conclusion
For the task of translating natural language into first-order logic, we have fined-tuned several large language models using the MALLS and Willow datasets, and have conducted experiments with several settings. An important aspect was the provision or learning of a list of predicate names that shall occur in the translation. While the provision of ground-truth predicate names greatly helps, learning them in a two-step process did not lead to good results. Our fine-tuned Flan-T5-XXL model with 11 billion parameters, can outperform baselines like GPT-4o, LogicLLaMA and even DeepSeek-R1-0528, as well as a symbolic baseline. The accuracy of translations of natural language arguments from the FOLIO dataset is comparable to the baseline (note that here only (in)validity of the translated argument is considered). The source code for our fine-tuned models is available on GitHub, and the models themselves can be accessed on Hugging Face. A common mistake in NL to FOL translation that is also present in the MALLS dataset is the confusion of hard-to-recognise Aristotelian forms: “A pineapple is a fruit with spiky skin” is translated to
@@
We observe that reasoning strategies such as CoT can sometimes degrade performance in NL-to-FOL translation, as initially correct outputs are overwritten during the reasoning process. This aligns with the illusion of thinking phenomenon described by Apple [illusion-of-thinking]. Future work may benefit from a two-stage approach, where base model outputs are first generated and then verified or refined using a reasoning model, as suggested by Thatikonda et al [thatikonda2024strategies].